Sketch the level curves f−1(c) for the following functions: (a) f(x,y) = x2 + y2, c = 0,1,2,3,4.
1
Expert's answer
2020-10-12T18:13:38-0400

Given, "f:\mathbb{R}^2\rightarrow \mathbb{R}" is a smooth map such that "f" is defined as below

"f(x,y)=x^2+y^2"

Now, when c=0 then


"f(x,y)=0\implies x^2+y^2=0\implies x=y=0"

Thus, "f^{-1}(c=0)=\{(0,0)\}" .

Now, for positive "c" ,


"f(x,y)=c=x^2+y^2"

Represents locus of radius "\sqrt{c}" and center is (0,0), thus


"f^{-1}(c)=\{(x,y)\in \mathbb{R}^2:f(x,y)=c\}"

is the set of points on the circle mentioned above.

Thus, level curves for c=0,1,2,3,4 will be



Where, origin is "f^{-1}(0)" , Blue, Green, Violet and black circles are respectively "f^{-1}(1),f^{-1}(2),f^{-1}(3)\&f^{-1}(4)" respectively.